OpenStudy (anonymous):

x^2+y^2+8x-8y-17=0 complete the square and write the equation in standard form.

OpenStudy (anonymous):

\[x^2+8x+y^2-8y=17\] \[(x+4)^2+(y-4)^2=17+16+16\] is a start

OpenStudy (anonymous):

actually that is all you have to do, just add on the right